Python))

  1. Data types of Python. Write the python script to show the different examples.

  1. Write the Python script to present ALL the mathematical arithmetic operations.

  1. Write the Python script to display your - first name, last name, ID, college name and the program of study.

  1. Write the Python script to receive the number of seconds as input and display the hours, minutes and seconds as output.

  1. Assume the variable x to be with the string of "Day by day, dear GOD, of thee three things I pray - to see thee more clearly; to love thee more dearly; to follow thee more nearly, Day by Day - St. Richard of ChiChester”.

Use the Python String Methods: str(), upper(), lower(), count(), find(), replace() & len(). Display the results.

  1. Write the program to determine how much to tip the server in a restaurant. The tip should be 15% of the bill, with a minimum of $2.
  2. A copy center charges 5 cents per copy for the first 100 copies and 3 cents per copy for each additional copy. Write a program that requests the number of copies as input and display the total cost.
  3. Federal law requires that hourly employees be paid “Time and a half” for work in excess of 40 hours in a week. For example, if a person’s hourly wage is $ 12 and he/ she works 60 hours in a week, the person’s gross pay should be:

             (40 * 12) + (1.5 * 12* (60 – 40)) = $840

  1. Display the table of numbers and the square.
  2. Develop the code to find the minimum, maximum and average of a sequence of non- negative numbers entered by the user. The user is told to enter the number -1 to indicate the end of data entry.  
Answer #1

1.

CODE

a = 5

print(a, "is of type", type(a))

a = 2.0

print(a, "is of type", type(a))

a = 1+2j

print(a, "is complex number?", isinstance(1+2j,complex))

a = [5,10,15,20,25,30,35,40]

# a[2] = 15

print("a[2] = ", a[2])

# a[0:3] = [5, 10, 15]

print("a[0:3] = ", a[0:3])

# a[5:] = [30, 35, 40]

print("a[5:] = ", a[5:])

t = (5,'program', 1+3j)

# t[1] = 'program'

print("t[1] = ", t[1])

# t[0:3] = (5, 'program', (1+3j))

print("t[0:3] = ", t[0:3])

s = 'Hello world!'

# s[4] = 'o'

print("s[4] = ", s[4])

# s[6:11] = 'world'

print("s[6:11] = ", s[6:11])

a = {5,2,3,1,4}

# printing set variable

print("a = ", a)

# data type of variable a

print(type(a))

d = {1:'value','key':2}

print(type(d))

print("d[1] = ", d[1]);

print("d['key'] = ", d['key']);

2.

CODE

a = 21

b = 10

c = 0

c = a + b

print("Line 1 - Value of c is ", c)

c = a - b

print("Line 2 - Value of c is ", c)

c = a * b

print("Line 3 - Value of c is ", c)

c = a / b

print("Line 4 - Value of c is ", c)

c = a % b

print("Line 5 - Value of c is ", c)

a = 2

b = 3

c = a**b

print("Line 6 - Value of c is ", c)

a = 10

b = 5

c = a//b

print("Line 7 - Value of c is ", c)

NOTE: As per Chegg policy, I am allowed to answer only 2 coding questions (including sub-parts) on a single post. Kindly post the remaining questions separately and I will try to answer them. Sorry for the inconvenience caused.